OpenStudy (anonymous):

What is the simplified form of the expression 3[14 ÷ (42 – 32) – 5]? (1 point)–9 27 21 6

OpenStudy (anonymous):

3(14÷10-5)= 3((7/5)-5)= 3(-18/5)=-54/5